# Spatial-Frequency Dual Domain Attention Network For Medical Image Segmentation thanks: The corresponding authors: Xueshuo Xie* and Tao Li*

### v1

@@ Abstract
In medical images, various types of lesions often manifest significant differences in their shape and texture. Accurate medical image segmentation demands deep learning models with robust capabilities in multi-scale and boundary feature learning. However, previous networks still have limitations in addressing the above issues. Firstly, previous networks simultaneously fuse multi-level features or employ deep supervision to enhance multi-scale learning. However, this may lead to feature redundancy and excessive computational overhead, which is not conducive to network training and clinical deployment. Secondly, the majority of medical image segmentation networks exclusively learn features in the spatial domain, disregarding the abundant global information in the frequency domain. This results in a bias towards low-frequency components, neglecting crucial high-frequency information. To address these problems, we introduce SF-UNet, a spatial-frequency dual-domain attention network. It comprises two main components: the Multi-scale Progressive Channel Attention (MPCA) block, which progressively extract multi-scale features across adjacent encoder layers, and the lightweight Frequency-Spatial Attention (FSA) block, with only 0.05M parameters, enabling concurrent learning of texture and boundary features from both spatial and frequency domains. We validate the effectiveness of the proposed SF-UNet on three public datasets. Experimental results show that compared to previous state-of-the-art (SOTA) medical image segmentation networks, SF-UNet achieves the best performance, and achieves up to 9.4% and 10.78% improvement in DSC and IOU. Codes will be released at https://github.com/nkicsl/SF-UNet .
@@ I Introduction
Medical image segmentation is the essential task and prerequisite for computer-aided diagnosis [1]. It refers to assigning each pixel a corresponding label to identify tissues or lesions. Deep learning-based medical image segmentation methods improve the segmentation accuracy, significantly reducing the workload of experts and eliminating the subjectivity of manual annotation. For example, segmenting lesions in fundus images and dermoscopic images enables early screening and intervention for diabetic retinopathy [2] and can enhance the diagnostic efficiency of melanoma and improve patient survival rates [3]. Accurate segmentation of teeth in CBCT images is also an important prerequisite for computer-aided diagnosis in oral medicine [4] [5]. These cases all highlight the significant role of deep learning in medical image segmentation.
Convolutional Neural Networks (CNN) have been widely used in computer vision tasks due to their powerful feature extraction capabilities. In medical image segmentation, UNet [6] and its variants [7] [8] are representative deep learning networks and have achieved remarkable performances. However, such CNN segmentation models still have some limitations: Firstly, due to the limitation of kernel size, CNN often have a limited receptive field, lacking the ability to learn global information and long-range relationships. Although some large kernel convolutional networks [9] [10] have been proposed recently, they often require more computational resources. Secondly, in medical image segmentation, the multi-scale problem is a significant challenge because different types of lesions often exhibit large differences in shape and volume [11]. CNN-based encoders often gradually downsample the resolution of feature maps, which may lead to the neglect of some small lesions, causing semantic loss and ultimately affecting the reliability of segmentation results. Some works use deep supervision [12] or fuse features from all levels of the encoder [13] to enhance the network’s multi-scale capabilities, but these may still lead to significant computational overhead and feature redundancy.
Figure 1: This figure illustrates the impact of high-frequency and low-frequency components on the image content from the ISIC2018 dataset. The top and bottom rows represent gradually removing the high-frequency and low-frequency features with different filter sizes of the image, respectively. Here n represents the width of the filter and the red slashed area represents the removed spectrum.
Transformer was originally proposed in natural language processing[14] and quickly applied to computer vision[15]. For medical images, global information and long-range relationship are crucial [16] [17]. Some pure transformer [18] [19] and hybrid transformer networks [20] [21] have achieved good results in medical image segmentation. However, transformers are not without their flaws. On one hand, global self-attention has quadratic computational complexity, leading to larger computational overhead and data requirements compared to CNN [22]. On the other hand, some researches have shown that global self-attention primarily focuses on the low-frequency components of the image, potentially neglecting high-frequency information to some extent [23]. Low-frequency components correspond to the content and texture of the image, while high-frequency components correspond to the edges and details [24]. As shown in Fig. 1, in the first row, we gradually extract the high-frequency information of the image, it can be observed that the image becomes increasingly blurred with the reduction of details. In the second row, as we gradually extract the low-frequency information from the image, we can observe that the texture of the image is removed, leaving only the components with larger gradients such as hair or boundaries. Therefore, to achieve accurate segmentation results, the network should consider both low-frequency and high-frequency information, ensuring the integrity of texture, boundaries and details.
@@
Our contributions are summarized as follows:
  • $\bullet$
    We design the MPCA and FSA block, the former can effectively learn multi-scale channel features, while the latter simultaneously conducts feature learning in both spatial and frequency domains, preserving the integrity of high-frequency information while retaining the ability to learn global information.
  • $\bullet$
    Based on MPCA and FSA, we built a U-shaped Spatial-Frequency Dual Domain Attention Network with strong capability to capture multi-scale features. Unlike Transformers, it ensures the network’s global feature learning ability from the frequency domain and preserves high-frequency detail features.
  • $\bullet$
    Extensive experiments were conducted on three public datasets to validate the effectiveness of the proposed network. The results show that SF-UNet can outperform previous SOTA networks. We also conducted ablation studies to further demonstrate the effectiveness of each proposed block.
@@ II-A Overall Structure
The overall pipeline of SF-UNet is illustrated in Fig 2. The encoder of SF-UNet consists of four encoder blocks followed by a bottleneck layer, following the implementation of VGG16 [25]. The encoder and decoder in SF-UNet are accompanied by MPCA and FSA blocks. Specifically, given an image $I\in\mathbb{R}^{H\times W\times C}$, H, W and C represent the height, width and channel numbers of the image, respectively. First, the output features of four encoder blocks and bottle layer are denoted as $F_{i}\in\mathbb{R}^{\frac{H}{2^{i-1}}\times\frac{W}{2^{i-1}}\times C_{i}}$, where $i\in\{1,2,3,4\}$, respectively. Then, for each level of MPCA block, $F_{i}$ and $F_{i+1}$ are used as inputs to accomplish cross-scale channel feature learning. The output of each MPCA block will go through a corresponding FSA block for spatial and frequency dual domain feature learning. Finally, a prediction head is used to get the final segmentation result $O$. Next, we elaborate the proposed MPCA and FSA blocks, respectively.
@@ II-B Multi-scale Progressive Channel Attention block
II-B1 Motivation
The multi-scale challenge in medical image segmentation is prominent. To ensure that each decoder block possesses cross-scale feature information and avoids feature redundancy, we propose the MPCA block. It progressively integrates feature maps from the current scale and the next scale to enhance the network’s multi-scale feature learning capability. The MPCA block fuse multi-scale features in two aspects. First, it generates channel attention maps from two adjacent scales simultaneously. Second, at the output of the MPCA block, we merge the weighted feature maps of the current scale and the next scale. This ensures that the outputs of the MPCA blocks include multi-scale fused features.
II-B2 Structure Detail
@@ II-C Frequency-Spatial Attention block
II-C1 Motivation
In medical imaging, the high-frequency information hold paramount importance for accurate diagnosis, as they encapsulate the precise delineation and boundaries of lesions. This underscores the need for meticulous consideration of high-frequency features inherent in medical images. Our objective is to equip the network with global feature learning capabilities, while conserving the high-frequency characteristics of the images. Based on this, we propose the FSA block. Unlike previous works [26, 27, 28] that solely rely on spatial domain processing, the FSA block concurrently conducts feature learning in both frequency and spatial domains. Specifically, the FSA block operates on the input feature maps by processing them with frequency and spatial branches. In the frequency branch, the spatial features are transformed into the frequency domain using 2D Discrete Fourier Transform (DFT), and the low-frequency and high-frequency components are separated. Subsequently, a learnable filter is applied to the low-frequency component for adaptive feature learning from a global perspective. The original high-frequency components are then reintegrated with the filtered low-frequency components, and the resultant spatial domain features are obtained using 2D Inverse Discrete Fourier Transform (IDFT). Meanwhile, in the spatial branch, spatial domain feature learning is facilitated using a spatial attention block adapted from [29]. Finally, the outputs from the two branches are fused to generate the output of the FSA block.
II-C2 Structure Detail
@@ III-A1 ISIC-2018[30]
The ISIC-2018 segmentation dataset developed by the International Skin Imaging Collaboration (ISIC), and it is employed for skin lesion segmentation. This dataset comprises 2594 training images, 100 validation images, and 1000 test images. Each image is captured using professional dermatoscopic equipment and annotated by experts. For our experiments, we use the official training, validation, and test sets provided by ISIC.
@@ III-C Experimental results
To comprehensively evaluate the performance of SF-UNet, we conducted qualitative and quantitative experiments on three datasets, comparing against seven previous SOTA networks: UNet [6], PSPNet [35], DeepLabV3+ [36], TransUNet [20], Swin-UNet [37], ACC-UNet [13], and MGFuseSeg [38].
@@ III-C1 Results on ISIC-2018
Dermoscopy images present typical challenges with multi-scale issue and blurry boundaries [39]. Skin diseases vary greatly in size and texture, demanding networks to capture both low-level and high-level features. Moreover, the boundaries between skin lesions and normal skin often lack clarity, presenting a significant challenge in lesion boundary delineation. Table I shows the quantitative results on the ISIC-2018 dataset. It is evident that SF-UNet outperformed previous SOTA networks in terms of DSC and IOU on the test set, highlighting the robust multi-scale feature learning and boundary feature learning capabilities of SF-UNet.
@@ III-C3 Results on NKUT
As mentioned in [32], the challenge of the NKUT dataset are the multi-scale and semantic confusion problems. This is because children’s mandibular wisdom teeth germs are often small and can be easily overlooked during the progressive downsampling. Additionally, determining the boundary between teeth and alveolar bone also poses a significant challenge. Addressing these concerns necessitates the network with robust multi-scale and edge feature learning capabilities. Table III presents the quantitative results on the NKUT dataset. From the results, SF-UNet achieved the best performance in terms of AB and average performance, and achieved the second best performance in terms of MWT and SM. For TransUNet, despite outperforming SF-UNet in the segmentation of MWT and SM, it has two limitations. Firstly, its parameters (92.23M) are significantly higher than SF-UNet. Secondly, its performance in segmenting AB is subpar, indicating inadequate multi-scale and boundary information learning capabilities. Its high IOU in MWT and SM may stem from aggressive classification of ambiguous boundaries between teeth and bones as teeth, compromising AB accuracy to enhance MWT and SM accuracy. Therefore, considering these aspects comprehensively, SF-UNet remains competitive.
@@ IV Conclusion
In this paper, we proposed SF-UNet for medical image segmentation, which consists of two blocks: MPCA and FSA blocks. We conducted extensive experiments on three public datasets, and the results showed that SF-UNet achieved the best performance, surpassing previous SOTA networks. Compared to other networks, SF-UNet demonstrated superior accuracy in segmenting textures and boundaries, as well as in distinguishing between various lesion types and sizes. Furthermore, we conducted ablation studies, providing quantitative results that validate the efficacy of all proposed blocks. In the future, we will continue to explore frequency-domain-based segmentation methods to achieve efficient, accurate, and computationally less intensive medical image segmentation, thereby establishing a robust framework for computer-aided diagnosis.

### v2

@@ Abstract
In medical images, various types of lesions often manifest significant differences in their shape and texture. The majority of medical image segmentation networks exclusively learn features in the spatial domain, disregarding the abundant global information in the frequency domain. This results in a bias towards low-frequency components, neglecting crucial high-frequency information. To address these problems, we introduce SF-UNet, a spatial-frequency dual-domain attention network. It comprises two main components: the Multi-scale Progressive Channel Attention (MPCA) block, which progressively extract multi-scale features across adjacent encoder layers, and the lightweight Frequency-Spatial Attention (FSA) block, with only 0.05M parameters, enabling concurrent learning of texture and boundary features from both spatial and frequency domains. We validate the effectiveness of the proposed SF-UNet on three public datasets. Experimental results show that compared to previous state-of-the-art (SOTA) medical image segmentation networks. Codes will be released at https://github.com/nkicsl/SF-UNet.
@@ I Introduction
Convolutional Neural Networks (CNN) have been widely used in computer vision tasks due to their powerful feature extraction capabilities. In medical image segmentation, UNet [1] and its variants [2] [3] are representative deep learning networks and have achieved remarkable performances. However, such CNN segmentation models still have some limitations: Firstly, due to the limitation of kernel size, CNN often have a limited receptive field, lacking the ability to learn global information and long-range relationships. Secondly, in medical image segmentation, the multi-scale problem is a significant challenge because different types of lesions often exhibit large differences in shape and volume [4]. CNN-based encoders often gradually downsample the resolution of feature maps, which may lead to the neglect of some small lesions, causing semantic loss and ultimately affecting the reliability of segmentation results.
Transformer was originally proposed in natural language processing[5] and quickly applied to computer vision[6]. For medical images, global information and long-range relationship are crucial [7]. However, transformers are not without their flaws. On one hand, global self-attention has quadratic computational complexity, leading to larger computational overhead and data requirements compared to CNN [8]. On the other hand, some researches have shown that global self-attention primarily focuses on the low-frequency components of the image, potentially neglecting high-frequency information to some extent [9].
@@ II-A Overall Structure
The overall pipeline of SF-UNet is illustrated in Fig 1. The encoder of SF-UNet consists of four encoder blocks followed by a bottleneck layer, following the implementation of VGG16 [10]. The encoder and decoder in SF-UNet are accompanied by MPCA and FSA blocks. Specifically, given an image $I\in\mathbb{R}^{H\times W\times C}$, H, W and C represent the height, width and channel numbers of the image, respectively. First, the output features of four encoder blocks and bottle layer are denoted as $F_{i}\in\mathbb{R}^{\frac{H}{2^{i-1}}\times\frac{W}{2^{i-1}}\times C_{i}}$, where $i\in\{1,2,3,4\}$, respectively. Then, for each level of MPCA block, $F_{i}$ and $F_{i+1}$ are used as inputs to accomplish cross-scale channel feature learning. The output of each MPCA block will go through a corresponding FSA block for spatial and frequency dual domain feature learning. Finally, a prediction head is used to get the final segmentation result $O$.
@@ III-A1 ISIC-2018[12]
This dataset comprises 2594 training images, 100 validation images, and 1000 test images. Each image is captured using professional dermatoscopic equipment and annotated by experts. For our experiments, we use the official training, validation, and test sets provided by ISIC.
@@ III-C1 Results on ISIC-2018
Table I shows the quantitative results on the ISIC-2018 dataset. It is evident that SF-UNet outperformed previous SOTA networks in terms of DSC and IOU on the test set, highlighting the robust multi-scale feature learning and boundary feature learning capabilities of SF-UNet.
@@ III-C3 Results on NKUT
Table III presents the quantitative results on the NKUT dataset. From the results, SF-UNet achieved the best performance in terms of AB and average performance, and achieved the second best performance in terms of MWT and SM. For TransUNet, despite outperforming SF-UNet in the segmentation of MWT and SM, it has two limitations. Firstly, its parameters (92.23M) are significantly higher than SF-UNet. Secondly, its performance in segmenting AB is subpar, indicating inadequate multi-scale and boundary information learning capabilities. Its high IOU in MWT and SM may stem from aggressive classification of ambiguous boundaries between teeth and bones as teeth, compromising AB accuracy to enhance MWT and SM accuracy. Therefore, considering these aspects comprehensively, SF-UNet remains competitive.
@@ IV Conclusion
In this paper, we proposed SF-UNet for medical image segmentation, which consists of two blocks: MPCA and FSA blocks. We conducted extensive experiments on three public datasets, and the results showed that SF-UNet achieved the best performance, surpassing previous SOTA networks. Compared to other networks, SF-UNet demonstrated superior accuracy in segmenting textures and boundaries, as well as in distinguishing between various lesion types and sizes.